OpenStudy (anonymous):

What is the 7th term of the geometric sequence 3, 9, 27

OpenStudy (anonymous):

\[3^{n}\] n=7 so..

OpenStudy (anonymous):

well.. if a_1=3 ... \[\LARGE q=\frac{9}{3}=3 \] \[\LARGE a_n=a_1\cdot q^{n-1}\] \[\LARGE a_7=3\cdot 3^{7-1}\] ...

OpenStudy (anonymous):

can you do it now ?

OpenStudy (anonymous):

yup i got it

OpenStudy (anonymous):

if u want to solved more geometri problem just do like @kreshnik said..